# Hetero-EUCLID: Interpretable model discovery for heterogeneous hyperelastic materials using stress-unsupervised learning11footnotemark: 1

### v1

@@ 3.2 Pattern case study
Figure 12: Model discovery for the multiple inner inclusions pattern (7(c)), with sub-figures (b)–(g) corresponding to the zero noise case. (a) displays the ground truth material segment IDs. (b)–(d) show the displacement fields in the $x$, $y$, and $z$ directions, respectively. (e) presents the residual force norm distribution (Equation 24) from the homogenized model. (f) illustrates the predicted material segments from the segmentation algorithm. (g) highlights the sub-sampled nodes from the $\mathcal{D}^{\text{free}}$ subset used to construct the heterogenized model, which is subsequently solved using the Bayesian framework.
Pattern Type: Multiple inner inclusions $|$ Mesh elements (C3D6): 10,816 Case Material ID Constitutive model $\theta_{1}$ (MPa) $\theta_{2}$ (MPa) $\theta_{3}$ (MPa) $\theta_{4}$ (MPa) $\theta_{5}$ (MPa) $\theta_{6}$ (MPa) Free nodes sub-sampling Ground Truth Refer (a) 1 ISH 4.00 0.50 0.30 - - 21.00 - 2 $\text{NH2}_{\text{a}}$ 5.40 - - - - 15.00 3 HW 1.00 0.15 - 0.02 0.00 10.00 4 $\text{NH2}_{\text{b}}$ 1.80 - - - - 6.00 Moderate Noise ($\sigma_{u}=2\times 10^{-3}$) 2 ISH 4.09 $\pm$ 0.07 0.51 $\pm$ 0.00 0.29 $\pm$ 0.02 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 21.12 $\pm$ 0.02 10% 5 $\text{NH2}_{\text{a}}$ 5.34 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 14.82 $\pm$ 0.00 3 HW 0.67 $\pm$ 0.07 0.52 $\pm$ 0.03 0.01 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 9.63 $\pm$ 0.14 4 $\text{NH2}_{\text{b}}$ 0.06 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 2.81 $\pm$ 0.00 High Noise ($\sigma_{u}=5\times 10^{-3}$) Refer (f) 2 ISH 4.91 $\pm$ 1.74 0.49 $\pm$ 0.06 0.17 $\pm$ 0.07 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 21.05 $\pm$ 0.13 10% 5 $\text{NH2}_{\text{a}}$ 4.91 $\pm$ 0.26 0.00 $\pm$ 0.03 0.00 $\pm$ 0.03 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 13.77 $\pm$ 0.52 3 HW 1.40 $\pm$ 5.76 0.62 $\pm$ 0.09 0.00 $\pm$ 0.00 0.01 $\pm$ 0.01 0.02 $\pm$ 0.06 6.10 $\pm$ 0.76 4 $\text{NH2}_{\text{b}}$ 0.68 $\pm$ 0.51 0.05 $\pm$ 0.05 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 3.16 $\pm$ 0.78
Table 7: True vs. predicted material parameters for multiple inner inclusions pattern with noise using Bayesian-EUCLID framework.
Figure 13: Model discovery for the multiple inner inclusions pattern (7(c)), with sub-figures (b)–(g) corresponding to the high noise case. (a) displays the ground truth material segment IDs. (b)–(d) show the displacement fields in the $x$, $y$, and $z$ directions, respectively. (e) presents the residual force norm distribution (Equation 24) from the homogenized model. (f) illustrates the predicted material segments from the segmentation algorithm. (g) highlights the sub-sampled nodes from the $\mathcal{D}^{\text{free}}$ subset used to construct the heterogenized model, which is subsequently solved using the Bayesian framework.
Pattern Type: Cahn-Hilliard $|$ Mesh elements (C3D6): 21,040 Case Material ID Constitutive model $\theta_{1}$ (MPa) $\theta_{2}$ (MPa) $\theta_{3}$ (MPa) $\theta_{4}$ (MPa) $\theta_{5}$ (MPa) $\theta_{6}$ (MPa) Free nodes sub-sampling Ground Truth, (a) 1 $\text{NH2}_{\text{a}}$ 5.40 - - - - 15.00 - 2 $\text{NH2}_{\text{b}}$ 1.80 - - - - 6.00 No Noise Refer (f) 5 $\text{NH2}_{\text{a}}$ 5.40 $\pm$ 0.01 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 15.05 $\pm$ 0.03 6 $\text{NH2}_{\text{b}}$ 1.78 $\pm$ 0.00 0.00 $\pm$ 0.03 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 5.94 $\pm$ 0.01 3 $\text{NH2}_{\text{a}}$ 5.42 $\pm$ 0.01 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 15.12 $\pm$ 0.03 5% 4 $\text{NH2}_{\text{b}}$ 1.80 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 6.00 $\pm$ 0.00 2 $\text{NH2}_{\text{a}}$ 5.42 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 15.02 $\pm$ 0.00
Table 8: True vs. predicted material parameters for the Cahn-Hilliard pattern without noise using the Bayesian-EUCLID framework.
Figure 14: Model discovery for the Cahn-Hilliard pattern (7(d)), with sub-figures (b)–(g) corresponding to the zero noise case. (a) displays the ground truth material segment IDs. (b)–(d) show the displacement fields in the $x$, $y$, and $z$ directions, respectively. (e) presents the residual force norm distribution (Equation 24) from the homogenized model. (f) illustrates the predicted material segments from the segmentation algorithm. (g) highlights the sub-sampled nodes from the $\mathcal{D}^{\text{free}}$ subset used to construct the heterogenized model, which is subsequently solved using the Bayesian framework.
Pattern Type: Cahn-Hilliard $|$ Mesh elements (C3D6): 21,040 Case Material ID Constitutive model $\theta_{1}$ (MPa) $\theta_{2}$ (MPa) $\theta_{3}$ (MPa) $\theta_{4}$ (MPa) $\theta_{5}$ (MPa) $\theta_{6}$ (MPa) Free nodes sub-sampling Ground Truth, (a) 1 $\text{NH2}_{\text{a}}$ 5.40 - - - - 15.00 - 2 $\text{NH2}_{\text{b}}$ 1.80 - - - - 6.00 Moderate Noise ($\sigma_{u}=2\times 10^{-3}$) 5 $\text{NH2}_{\text{a}}$ 6.20 $\pm$ 0.09 0.00 $\pm$ 0.00 0.05 $\pm$ 0.05 0.00 $\pm$ 0.00 0.01 $\pm$ 0.01 16.76 $\pm$ 0.16 6 $\text{NH2}_{\text{b}}$ 1.82 $\pm$ 0.02 0.00 $\pm$ 0.03 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 6.03 $\pm$ 0.07 3 $\text{NH2}_{\text{a}}$ 5.28 $\pm$ 0.05 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 14.64 $\pm$ 0.12 5% 4 $\text{NH2}_{\text{b}}$ 1.85 $\pm$ 0.01 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 6.12 $\pm$ 0.04 2 $\text{NH2}_{\text{a}}$ 4.44 $\pm$ 0.08 0.46 $\pm$ 0.02 0.07 $\pm$ 0.02 0.01 $\pm$ 0.01 0.00 $\pm$ 0.00 15.13 $\pm$ 0.07 High Noise ($\sigma_{u}=5\times 10^{-3}$) Refer (f) 5 $\text{NH2}_{\text{a}}$ 6.74 $\pm$ 0.06 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 17.84 $\pm$ 0.20 6 $\text{NH2}_{\text{b}}$ 1.72 $\pm$ 0.03 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 5.71 $\pm$ 0.09 3 $\text{NH2}_{\text{a}}$ 5.03 $\pm$ 0.04 0.00 $\pm$ 0.00 0.03 $\pm$ 0.01 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 14.18 $\pm$ 0.11 10% 4 $\text{NH2}_{\text{b}}$ 1.66 $\pm$ 0.02 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 5.52 $\pm$ 0.06 2 $\text{NH2}_{\text{a}}$ 5.39 $\pm$ 0.10 0.02 $\pm$ 0.03 0.02 $\pm$ 0.02 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 16.19 $\pm$ 0.11
Table 9: True vs. predicted material parameters for the Cahn-Hilliard pattern with noise using the Bayesian-EUCLID framework.
Figure 15: Model discovery for the Cahn-Hilliard pattern (7(d)), with sub-figures (b)–(g) corresponding to the high noise case. (a) displays the ground truth material segment IDs. (b)–(d) show the displacement fields in the $x$, $y$, and $z$ directions, respectively. (e) presents the residual force norm distribution (Equation 24) from the homogenized model. (f) illustrates the predicted material segments from the segmentation algorithm. (g) highlights the sub-sampled nodes from the $\mathcal{D}^{\text{free}}$ subset used to construct the heterogenized model, which is subsequently solved using the Bayesian framework.
@@
Using the homogenized model formulation, a distinct imbalance in the residual force norm is observed along the complex Cahn–Hilliard pattern boundaries, as shown in Figure 14e, 15e, and S4e. Material region segmentation is performed using the threshold parameter $\lambda=1.5$, resulting in the prediction of six material segments. The segmentation output is visualized in Figure 14f, 15f, and S4f, where segment ID 1 contains a small number of misclustered elements under both noise-free and noisy conditions. The remaining segments (IDs 2-6) exhibit well-defined and accurate material interfaces. Subsequently, the material parameter identification is carried out using 5% sub-sampling of the free nodes, and the noise-free results are reported in Table 8. Relative to the ground truth, material segments with IDs 2, 3, and 5 recover $\text{NH2}_{\text{a}}$ model coefficients $\theta_{1}$ and $\theta_{6}$ close to 5.40 MPa and 15.00 MPa, respectively.
@@
Similarly, the corresponding predictions of $\theta_{1}$ and $\theta_{6}$ for material segments 4 and 6 closely match their $\text{NH2}_{\text{b}}$ ground truth values of 1.80 MPa and 6.00 MPa, respectively. All remaining material parameters $\theta_{2-5}$ are correctly predicted as zero with low standard deviation. A key advantage of the Hetero-EUCLID framework lies in its interpretability. While the segmentation algorithm identifies five major material regions, the predicted model parameters provide meaningful physical insight into material similarity. It can be inferred that segments 2, 3, and 5 exhibit nearly identical Neo-Hookean model coefficients, indicating that these segments represent the same material. Similarly, material regions with IDs 4 and 6 are identified as belonging to a second distinct material type. Thus, Hetero-EUCLID segments the domain and enables grouping regions based on interpretable material parameters, effectively identifying five spatial zones representing two unique material classes.
@@
Table 9 summarizes a comparative study of material parameter prediction performance for the Cahn-Hilliard pattern specimen under moderate and high noise settings. The resulting predictions of $\theta_{1}$ and $\theta_{6}$ across all identified Neo-Hookean regions lie within an acceptable tolerance of the ground truth and are associated with low standard deviations, indicating high confidence. Additionally, the remaining coefficients $\theta_{2-5}$ are correctly predicted as zero, promoting sparsity. Supplementary Figure S8 demonstrates strong agreement between the strain energy density plots computed from the predicted model parameters and those from the corresponding ground truth. These results demonstrate the robustness and interpretability of the Hetero-EUCLID framework for complex heterogeneous configurations under high noise conditions.
@@ 3.3 Non-native mesh
In subsection 3.2, we evaluated the performance of the Hetero-EUCLID framework across various case studies. For all configurations, the computational domain was discretized using wedge elements based on a known element connectivity matrix. Notably, the same mesh (referred to as the native mesh) used in the forward FEM simulations for generating displacement data was also used for the inverse formulation. In this section, we investigate further to emulate DIC experimental conditions more pragmatically, by evaluating the Hetero-EUCLID framework on a non-native mesh. A non-native mesh refers to a discretization where the element edges do not necessarily align with the underlying material interface boundaries. To investigate this scenario, we use the complex Type 4 Cahn–Hilliard heterogeneous pattern and discretize the domain using a uniform grid and generate wedge elements using Delaunay triangulation [59] on the surface, resulting in 19,602 uniformly distributed wedge elements. Figure 16a provides a comparative visualization between the native (I) and non-native (II) meshes, showing a zoomed-in view near a material interface. It can be observed that the native mesh aligns precisely with the material boundary, while the non-native mesh does not. In DIC experiments, the material interface geometry is inherently unknown during meshing; thus, this non-native mesh study is key to evaluating the robustness and practical applicability of Hetero-EUCLID framework under realistic experimental conditions.
Using the displacement field generated from ABAQUS/CAE for the Cahn–Hilliard pattern under non-equi-biaxial loading, we perform FE-based interpolation to compute full-field displacements at the nodes of the non-native mesh. As the FE-based interpolation is expected to have inherent noise, we do not add further Gaussian noise to the displacement data as was done in previous cases. Figure 16b illustrates the distribution of the residual force norm, where high magnitudes highlight the material interface.
Pattern Type: Cahn-Hilliard $|$ Non-native mesh elements (C3D6): 19,602 Case Material ID Constitutive model $\theta_{1}$ (MPa) $\theta_{2}$ (MPa) $\theta_{3}$ (MPa) $\theta_{4}$ (MPa) $\theta_{5}$ (MPa) $\theta_{6}$ (MPa) Free nodes sub-sampling Ground Truth, (a) 1 $\text{NH2}_{\text{a}}$ 5.40 - - - - 15.00 - 2 $\text{NH2}_{\text{b}}$ 1.80 - - - - 6.00 Using Ordinary Least Squares Regression 3 $\text{NH2}_{\text{a}}$ 7.58 0.29 0.74 -0.51 -0.30 21.14 5 $\text{NH2}_{\text{b}}$ 2.03 -0.15 -0.07 0.01 0.00 4.22 2 $\text{NH2}_{\text{a}}$ 4.95 -0.36 -0.19 0.13 0.01 12.12 - 4 $\text{NH2}_{\text{b}}$ 2.25 -0.24 -0.07 0.01 0.00 3.68 6 $\text{NH2}_{\text{a}}$ 8.98 -1.65 -0.49 0.58 -0.06 18.88 Using Bayesian framework Refer (c) 3 $\text{NH2}_{\text{a}}$ 5.94 $\pm$ 0.27 0.11 $\pm$ 0.12 0.00 $\pm$ 0.00 0.01 $\pm$ 0.04 0.00 $\pm$ 0.00 17.20 $\pm$ 0.61 5 $\text{NH2}_{\text{b}}$ 1.31 $\pm$ 0.03 0.02 $\pm$ 0.01 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 4.39 $\pm$ 0.06 2 $\text{NH2}_{\text{a}}$ 4.84 $\pm$ 0.25 0.20 $\pm$ 0.03 0.01 $\pm$ 0.01 0.01 $\pm$ 0.00 0.00 $\pm$ 0.00 14.26 $\pm$ 0.69 10% 4 $\text{NH2}_{\text{b}}$ 1.76 $\pm$ 0.01 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 5.64 $\pm$ 0.02 6 $\text{NH2}_{\text{a}}$ 5.89 $\pm$ 0.04 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.01 15.26 $\pm$ 0.02
Table 10: True vs. predicted material parameters for the Cahn-Hilliard pattern using non-native meshing.
@@
Figure 16: Model discovery for the Cahn–Hilliard pattern (7(d)) using non-native meshing. (a) shows the ground truth material segment IDs with visual comparison of (I) native and (II) non-native mesh views. (b) presents the residual force norm distribution (Equation 24) from the homogenized model. (c) illustrates the predicted material segments obtained through the segmentation algorithm. Table 10 summarizes the ground truth and predicted material parameters obtained by solving the heterogenized model using the ordinary least squares regression and Bayesian framework.
@@
Subsequently, the segmentation is performed using the random seed-based island growth algorithm, and the resulting material segments are shown in Figure 16c. While the algorithm successfully identifies five major material zones, the interfaces appear less sharp, which is expected and justified given the non-aligned structure of the non-native mesh grid compared to those observed with the native mesh configuration.

### v2

@@ 3.3 Pattern case study
Figure 10: Model discovery for the multiple inner inclusions pattern (7(c)). The ground truth material segment IDs are shown in (a), and the displacement fields in the $x$, $y$, and $z$ directions are provided in (b)–(d). For the zero-noise case, (e) depicts the residual force norm distribution (Equation 26) from the homogenized model, (f) illustrates the material segments identified through the segmentation algorithm, and (g) highlights the sub-sampled nodes from the $\mathcal{D}^{\text{free}}$ set used to construct the heterogenized model. The high-noise case follows a similar structure in panels (h)–(j), showing how the residual force field, segmentation, and sub-sampling behave under high noise levels. In both scenarios, the final heterogenized model is solved using the Bayesian framework to recover the region-wise constitutive parameters as listed in Table 4.
Pattern Type 4: Cahn-Hilliard pattern
In this type of domain heterogeneity, we investigate the Cahn-Hilliard pattern, which is often used to represent the biological pattern formation in soft tissue mechanics. 7(d) illustrates this configuration, featuring five adjacent material regions alternately composed of two Neo-Hookean ($\text{NH2}_{\text{a}},\text{NH2}_{\text{b}}$) constitutive models. This specimen serves to evaluate the ability of Hetero-EUCLID to distinguish material sub-regions when multiple zones share identical mechanical properties. The domain is meshed with 21,040 uniformly distributed wedge-type elements. Using ABAQUS/CAE, we generate the full-field computational displacement dataset for quasi-static non-equi-biaxial loading as shown in Figure 11b-d. The Cahn–Hilliard heterogeneous configuration is analyzed under noise-free, high, and moderate-noise conditions, with the corresponding results summarized in Figure 11e-g, 11h-j, and Supplementary Figure S4e-g, respectively.
@@
Using the homogenized model formulation, a distinct imbalance in the residual force norm is observed along the complex Cahn–Hilliard pattern boundaries, as shown in Figure 11e, 11h, and S4e. Material region segmentation is performed using the threshold parameter $\lambda=1.5$, resulting in the prediction of six material segments. The segmentation output is visualized in Figure 11f, 11i, and S4f, where segment ID 1 contains a small number of misclustered elements under both noise-free and noisy conditions. The remaining segments (IDs 2-6) exhibit well-defined and accurate material interfaces. Subsequently, the material parameter identification is carried out using 5% sub-sampling of the free nodes, and the noise-free results are reported in Table 5. Relative to the ground truth, material segments with IDs 2, 3, and 5 recover $\text{NH2}_{\text{a}}$ model coefficients $\theta_{1}$ and $\theta_{19}$ close to 5.40 MPa and 15.00 MPa, respectively.
Pattern Type 4: Cahn-Hilliard $|$ Mesh elements (C3D6): 21,040 $|$ Activated feature library terms, $\boldsymbol{n_{\mathit{f}}=6}$ Case Material ID Constitutive model $\theta_{1}$ (MPa) $\theta_{2}$ (MPa) $\theta_{3}$ (MPa) $\theta_{4}$ (MPa) $\theta_{5}$ (MPa) $\theta_{19}$ (MPa) Free nodes sub-sampling Ground Truth Refer 11(a) 1 $\text{NH2}_{\text{a}}$ 5.40 - - - - 15.00 - 2 $\text{NH2}_{\text{b}}$ 1.80 - - - - 6.00 No Noise ($\sigma_{u}=0$) Refer 11(f) 5 $\text{NH2}_{\text{a}}$ 5.44 $\pm$ 0.01 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 15.08 $\pm$ 0.06 6 $\text{NH2}_{\text{b}}$ 1.79 $\pm$ 0.01 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 5.97 $\pm$ 0.02 3 $\text{NH2}_{\text{a}}$ 5.41 $\pm$ 0.03 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 15.04 $\pm$ 0.05 5% 4 $\text{NH2}_{\text{b}}$ 1.80 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 5.99 $\pm$ 0.01 2 $\text{NH2}_{\text{a}}$ 5.43 $\pm$ 0.01 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 15.03 $\pm$ 0.01 Moderate Noise ($\sigma_{u}=2\times 10^{-3}$) Refer S4(f) 5 $\text{NH2}_{\text{a}}$ 6.36 $\pm$ 0.03 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 16.88 $\pm$ 0.11 6 $\text{NH2}_{\text{b}}$ 1.84 $\pm$ 0.01 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 6.07 $\pm$ 0.04 3 $\text{NH2}_{\text{a}}$ 5.31 $\pm$ 0.03 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 14.67 $\pm$ 0.08 5% 4 $\text{NH2}_{\text{b}}$ 1.78 $\pm$ 0.01 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 5.93 $\pm$ 0.03 2 $\text{NH2}_{\text{a}}$ 4.76 $\pm$ 0.05 0.37 $\pm$ 0.01 0.07 $\pm$ 0.01 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 15.48 $\pm$ 0.05 High Noise ($\sigma_{u}=5\times 10^{-3}$) Refer 11(i) 5 $\text{NH2}_{\text{a}}$ 6.74 $\pm$ 0.22 0.00 $\pm$ 0.00 0.10 $\pm$ 0.10 0.00 $\pm$ 0.00 0.00 $\pm$ 0.01 18.32 $\pm$ 0.23 6 $\text{NH2}_{\text{b}}$ 1.78 $\pm$ 0.03 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 5.86 $\pm$ 0.08 3 $\text{NH2}_{\text{a}}$ 5.03 $\pm$ 0.06 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 13.81 $\pm$ 0.16 10% 4 $\text{NH2}_{\text{b}}$ 1.67 $\pm$ 0.02 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 5.51 $\pm$ 0.07 2 $\text{NH2}_{\text{a}}$ 5.45 $\pm$ 0.11 0.17 $\pm$ 0.04 0.02 $\pm$ 0.02 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 16.28 $\pm$ 0.12
Table 5: True vs. predicted material parameters for the Cahn-Hilliard pattern (7(d)) across different noise levels using the Bayesian-EUCLID framework.
@@
Similarly, the corresponding predictions of $\theta_{1}$ and $\theta_{19}$ for material segments 4 and 6 closely match their $\text{NH2}_{\text{b}}$ ground truth values of 1.80 MPa and 6.00 MPa, respectively. All remaining material parameters $\theta_{2-5}$ are correctly predicted as zero with low standard deviation. A key advantage of the Hetero-EUCLID framework lies in its interpretability. While the segmentation algorithm identifies five major material regions, the predicted model parameters provide meaningful physical insight into material similarity. It can be inferred that segments 2, 3, and 5 exhibit nearly identical Neo-Hookean model coefficients, indicating that these segments represent the same material. Similarly, material regions with IDs 4 and 6 are identified as belonging to a second distinct material type. Thus, Hetero-EUCLID segments the domain and enables grouping of regions based on interpretable material parameters, effectively identifying five spatial zones that represent two unique material classes.
Figure 11: Model discovery for the Cahn-Hilliard pattern (7(d)). The ground truth material segment IDs are shown in (a), and the displacement fields in the $x$, $y$, and $z$ directions are provided in (b)–(d). For the zero-noise case, (e) depicts the residual force norm distribution (Equation 26) from the homogenized model, (f) illustrates the material segments identified through the segmentation algorithm, and (g) highlights the sub-sampled nodes from the $\mathcal{D}^{\text{free}}$ set used to construct the heterogenized model. The high-noise case follows a similar structure in panels (h)–(j), showing how the residual force field, segmentation, and sub-sampling behave under high noise levels. In both scenarios, the final heterogenized model is solved using the Bayesian framework to recover the region-wise constitutive parameters as listed in Table 5.
@@
Table 5 summarizes a comparative study of material parameter prediction performance for the Cahn-Hilliard pattern specimen under moderate and high noise settings. The resulting predictions of $\theta_{1}$ and $\theta_{19}$ across all identified Neo-Hookean regions lie within an acceptable tolerance of the ground truth and are associated with low standard deviations, indicating high confidence. Additionally, the remaining coefficients $\theta_{2-5}$ are correctly predicted as zero, promoting sparsity. Supplementary Figure S10 demonstrates strong agreement between the strain energy density plots computed from the predicted model parameters and those from the corresponding ground truth. These results demonstrate the robustness and interpretability of the Hetero-EUCLID framework for complex heterogeneous configurations under high noise conditions.
@@ 3.4 Non-native mesh
Figure 12: Model discovery for the Cahn-Hilliard pattern (7(d)) using non-native meshing. (a) shows the ground truth material segment IDs with visual comparison of (I) native and (II) non-native mesh views. For the zero-noise case, (b) depicts the residual force norm distribution (Equation 26) from the homogenized model, and (c) illustrates the material segments identified through the segmentation algorithm. The high-noise case follows a similar structure in panels (d)–(e), showing how the residual force field and segmentation behave under high noise levels. In both scenarios, the final heterogenized model is solved using the Bayesian framework to recover the region-wise constitutive parameters as listed in Table 6.
In subsection 3.3, we evaluated the performance of the Hetero-EUCLID framework across various case studies. For all configurations, the computational domain was discretized using wedge elements based on a known element connectivity matrix. Notably, the same mesh (referred to as the native mesh) used in the forward FEM simulations for generating displacement data was also used for the inverse formulation. In this section, we investigate further to emulate DIC experimental conditions more pragmatically, by evaluating the Hetero-EUCLID framework on a non-native mesh. A non-native mesh refers to a discretization where the element edges do not necessarily align with the underlying material interface boundaries. To investigate this scenario, we use the complex Type 4 Cahn–Hilliard heterogeneous pattern and discretize the domain using a uniform grid and generate wedge elements using Delaunay triangulation [Lee and Schachter, 1980] on the surface, resulting in 19,602 uniformly distributed wedge elements. Figure 12a provides a comparative visualization between the native (I) and non-native (II) meshes, showing a zoomed-in view near a material interface. It can be observed that the native mesh aligns precisely with the material boundary, while the non-native mesh does not. In DIC experiments, the material interface geometry is inherently unknown during meshing; therefore, this non-native mesh study is crucial for evaluating the robustness and practical applicability of the Hetero-EUCLID framework under realistic experimental conditions.
@@
Using the displacement field generated from ABAQUS/CAE for the Cahn–Hilliard pattern under non-equi-biaxial loading, we perform FE-based interpolation to compute full-field displacements at the nodes of the non-native mesh. Such an FE-based interpolation is expected to add noise to the displacement field. In this context, we evaluate the performance of our Hetero-EUCLID framework for two cases– one without any additional displacement noise (apart from the inherent interpolation noise) and another with additional, high, uncorrelated displacement noise $(\sigma_{u,\text{high}}=5\times 10^{-3})$. Subsequently, the segmentation is performed using the random seed-based island growth algorithm, and the resulting material segments are shown in Figure 12c and 12e. While the algorithm successfully identifies five major material zones, the interfaces appear less sharp, which is expected and justified given the non-aligned structure of the non-native mesh grid compared to those observed with the native mesh configuration.
Pattern Type 4: Cahn-Hilliard $|$ Non-native Mesh elements (C3D6): 19,602 $|$ Activated feature library terms, $\boldsymbol{n_{\mathit{f}}=6}$ Case Material ID Constitutive model $\theta_{1}$ (MPa) $\theta_{2}$ (MPa) $\theta_{3}$ (MPa) $\theta_{4}$ (MPa) $\theta_{5}$ (MPa) $\theta_{19}$ (MPa) Free nodes sub-sampling Ground Truth, 12(a) 1 $\text{NH2}_{\text{a}}$ 5.40 - - - - 15.00 - 2 $\text{NH2}_{\text{b}}$ 1.80 - - - - 6.00 No Noise ($\sigma_{u}=0$) Bayesian framework Refer 12(c) 6 $\text{NH2}_{\text{a}}$ 6.05 $\pm$ 0.96 0.12 $\pm$ 0.43 0.01 $\pm$ 0.01 0.00 $\pm$ 0.01 0.00 $\pm$ 0.00 17.00 $\pm$ 1.62 5 $\text{NH2}_{\text{b}}$ 1.29 $\pm$ 0.09 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 4.12 $\pm$ 0.27 3 $\text{NH2}_{\text{a}}$ 4.97 $\pm$ 0.99 0.12 $\pm$ 0.08 0.01 $\pm$ 0.03 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 14.19 $\pm$ 2.50 10% 4 $\text{NH2}_{\text{b}}$ 1.77 $\pm$ 0.01 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 5.61 $\pm$ 0.03 2 $\text{NH2}_{\text{a}}$ 5.89 $\pm$ 0.09 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.02 0.00 $\pm$ 0.00 15.25 $\pm$ 0.04 High Noise ($\sigma_{u}=5\times 10^{-3}$) (Ordinary Least Squares Regression) 5 $\text{NH2}_{\text{a}}$ 12.84 -2.99 -3.09 1.35 0.41 18.76 3 $\text{NH2}_{\text{b}}$ 1.31 0.24 -0.06 0.06 0.00 3.68 2 $\text{NH2}_{\text{a}}$ 1.86 1.05 0.71 -0.13 -0.07 9.59 - 4 $\text{NH2}_{\text{b}}$ 1.71 -0.02 -0.06 0.08 0.00 3.50 6 $\text{NH2}_{\text{a}}$ 8.38 -2.31 -1.92 1.94 -0.02 13.65 High Noise ($\sigma_{u}=5\times 10^{-3}$) Bayesian framework Refer 12(e) 5 $\text{NH2}_{\text{a}}$ 7.69 $\pm$ 0.07 0.00 $\pm$ 0.00 0.04 $\pm$ 0.04 0.00 $\pm$ 0.00 0.07 $\pm$ 0.01 18.82 $\pm$ 0.16 3 $\text{NH2}_{\text{b}}$ 1.55 $\pm$ 0.04 0.12 $\pm$ 0.01 0.00 $\pm$ 0.00 0.03 $\pm$ 0.00 0.00 $\pm$ 0.00 3.49 $\pm$ 0.04 2 $\text{NH2}_{\text{a}}$ 3.36 $\pm$ 0.11 0.42 $\pm$ 0.06 0.26 $\pm$ 0.03 0.01 $\pm$ 0.02 0.00 $\pm$ 0.00 10.27 $\pm$ 0.10 10% 4 $\text{NH2}_{\text{b}}$ 1.74 $\pm$ 0.02 0.00 $\pm$ 0.01 0.00 $\pm$ 0.00 0.04 $\pm$ 0.00 0.00 $\pm$ 0.00 3.46 $\pm$ 0.04 6 $\text{NH2}_{\text{a}}$ 5.66 $\pm$ 0.07 0.00 $\pm$ 0.00 0.13 $\pm$ 0.04 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00 13.21 $\pm$ 0.09
Table 6: True vs. predicted material parameters for the non-natively meshed Cahn-Hilliard pattern. The results are presented for three cases, (i) no additional displacement noise, solved using Bayesian regression framework, (ii) High uncorrelated displacement noise ($\sigma_{u,\mathrm{high}}=5\times 10^{-3}$) solved using the OLS regression approach, and (iii) High uncorrelated displacement noise ($\sigma_{u,\mathrm{high}}=5\times 10^{-3}$) solved using the Bayesian regression approach.